# LLM Agents for Education: Advances and Applications

### v1

@@
Abstract.
Large Language Model (LLM) agents have demonstrated remarkable capabilities in automating tasks and driving innovation across diverse educational applications. In this survey, we provide a systematic review of state-of-the-art research on LLM agents in education, categorizing them into two broad classes: (1) Pedagogical Agents, which focus on automating complex pedagogical tasks to support both teachers and students; and (2) Domain-Specific Educational Agents, which are tailored for specialized fields such as science education, language learning, and professional development. We comprehensively examine the technological advancements underlying these LLM agents, including key datasets, benchmarks, and algorithmic frameworks that drive their effectiveness. Furthermore, we discuss critical challenges such as privacy, bias and fairness concerns, hallucination mitigation, and integration with existing educational ecosystems. This survey aims to provide a comprehensive technological overview of LLM agents for education, fostering further research and collaboration to enhance their impact for the greater good of learners and educators alike.
Large language models, Adaptive learning, AI for education
*Equal contribution. Corresponding author.
copyright: nonejournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
@@
Artificial intelligence (AI) techniques are increasingly used in education to enable personalized learning and intelligent tutoring (Chen et al., 2020; Zhai et al., 2021; Pedro et al., 2019). While traditional educational data mining approaches (Shafiq et al., 2022; Khan and Ghosh, 2021; Abdelrahman et al., 2023; Song et al., 2022; Wang et al., 2022; Gao et al., 2021), such as knowledge tracing and cognitive diagnosis, have made significant progress in reshaping the human-learning paradigm by analyzing student behaviors and assessing knowledge states, they still face major challenges in real-world applications. These challenges include shallow contextual understanding, limited interactive capabilities, and difficulties in generating adaptive, personalized learning materials, etc (Zhu et al., 2024; Laak and Aru, 2024; Tan et al., 2023).
@@
The strong natural language understanding of Large Language Models (LLMs) and the task automation capabilities of LLM agents make them valuable for addressing challenges in education (Weng, 2023; Wang et al., 2024d). First, memory enables LLM agents to retain both long-term knowledge about students’ study habits and short-term context from real-time interactions, enhancing contextual understanding and ensuring personalized learning experiences across various educational tasks (Zhang et al., 2024a). Second, tool use allows LLM agents to access external resources, perform complex calculations, and retrieve real-time information, enabling them to automate intricate educational tasks such as grading, knowledge retrieval, and adaptive content generation, thereby overcoming limited interactivity and enhancing engagement (Gao et al., 2023). Third, planning supports structured learning by decomposing complex topics, predicting optimal learning paths, and dynamically adjusting instructional strategies, allowing LLM agents to autonomously guide students through personalized learning experiences (Huang et al., 2024a). By integrating these capabilities, LLM agents not only enhance understanding, engagement, and personalization but also automate complex educational workflows, making learning more adaptive, efficient, and scalable.
In this survey, we provide a comprehensive review of LLM agents in education. We begin by introducing LLM agents, followed by a general discussion of their potential applications in educational systems. Next, we categorize them into two broad classes: (1) Pedagogical Agents, which automate complex pedagogical tasks such as classroom simulation and learning resource recommendation for teachers. Also, they transform student support paradigms through applications like knowledge tracing, adaptive learning, and error detection and correction. (2) Domain-Specific Educational Agents, which are designed to address specialized challenges in different fields, including science education, language learning, and professional development. Furthermore, we discuss the key challenges in LLM agents for education, which need to be addressed to advance research and enable effective deployment in real-world educational applications. Finally, we compile comprehensive datasets, benchmarks, and evaluation methodologies to support and encourage further research involvement in the area of LLM agents for education. We summarize our contributions as follows:
Figure 1. The overview of LLM Agents for education.
  • Comprehensive review of LLM Agents in education. We analyze recent advancements in LLM-driven educational agents, discussing their applications, effectiveness, and challenges in automating pedagogical tasks and enhancing learning experiences.
  • Novel task-centric taxonomy. We propose a structured classification of LLM agents, categorizing them into Pedagogical Agents and Domain-Specific Educational Agents, providing a framework for understanding their roles and capabilities.
@@
  • Current challenges and future research directions. We analyze critical challenges that need to be addressed for the effective deployment of LLM agents in education, including issues related to privacy, bias and fairness concerns, hallucination, and integration into real-world educational ecosystems.
@@
  • Compilation of essential resources. We compile comprehensive datasets and benchmarks to support future research efforts and facilitate the development of more robust and effective LLM-driven educational solutions [^1] Due to the page limit, we present more details in Appendix A. .
2. LLM Agents for Education
LLM agents, powered by LLMs (OpenAI, 2022, 2023; Touvron et al., 2023), combine core components like memory, tool use, and planning to enhance the performance for diverse tasks (Xi et al., 2025; Durante et al., 2024; Tran et al., 2025; Liu et al., 2025; Yan et al., 2024d). They leverage long-term memory for foundational knowledge and short-term memory for real-time adaptation, enabling context-aware interactions. Active memory management, such as summarization and retrieval (Chen et al., 2023; Liang et al., 2023), helps maintain relevant information. To overcome LLM limitations, such as knowledge cutoff and calculation inefficiency (Liska et al., 2022), LLM agents integrate external tools like search engines, calculators, and databases (Qin et al., 2023; Zhuang et al., 2023), enabling real-time data access and multi-step tasks. Additionally, planning capabilities allow agents to decompose goals, simulate state transitions, and optimize actions for complex problem-solving (Yao et al., 2023; Valmeekam et al., 2023; Xie et al., 2024c).
Education, a field inherently focused on individualized learning and problem-solving, is one of the ideal scenarios for deploying LLM agents. These agents can significantly enhance the learning experience by providing personalized support tailored to the unique needs of each student. To achieve this, as illustrated in Figure 1, LLM agents must integrate several key capabilities, including memory, tool-use, and planning modules. The memory module enables the agent to retain an understanding of fundamental concepts, track the student’s past interactions and progress, and store personalized profiles to provide accurate responses. The tool-use module allows the agent to access specialized educational tools, such as knowledge tracing and recommendation models, as well as external resources like search engines, thereby expanding its functional capabilities and ensuring up-to-date, accurate information. The planning module empowers the agent to deconstruct the student’s objectives, plan tailored learning paths, and offer personalized feedback based on the student’s progress and performance. Together, these modules allow LLM agents to function as dynamic, responsive tutors capable of guiding students through both simple and complex tasks. By offering context-aware feedback, adapting to evolving needs, and tailoring learning experiences, LLM agents have the potential to revolutionize education. Their ability to provide scalable, adaptive, and efficient support can transform personalized education, making it more accessible, engaging, and effective. As shown in Figure 2, we categorize LLM agents for education into two types: Pedagogical Agents and Domain-Specific Educational Agents, and outline the roadmap for discussion in the following sections.
@@ 2. LLM Agents for Education
for tree= forked edges, grow’=0, calign=child edge, calign child=(n_children()+1)/2, , [LLM Agents for Education, root [Pedagogical Agents, sa [Teaching Assistance §3.1, ta [Classroom Simulation (CS), ta [EduAgent (Xu et al., 2024a); TeachTune (Jin et al., 2024b); CGMI (Jinxin et al., 2023) Classroom Simulacra (Xu et al., 2025);  Li et al. (2025a);  Bhowmik et al. (2024);  Zheng et al. (2025);  Hu et al. (2025b), ta_work ] ] [Feedback Comment Generation (FCG), ta [ PROF (Nair et al., 2024); SEFL (Zhang et al., 2025a); FreeText (Matelsky et al., 2023); AAAR-1.0 (Lou et al., 2024);  Guo et al. (2024a);  Estévez-Ayres et al. (2024);  Du et al. (2024) ,ta_work ] ] [Learning Resource
Recommendation (LRR), ta [Zaiane (2002); SKarREC (Li et al., 2024d);  Moon et al. (2024);  Abu-Rasheed et al. (2024b);  Abu-Rasheed et al. (2024a) ,ta_work ] ] ] [Student Support §3.2, ss [Adaptive Learning (AL), ss [EduAgent (Xu et al., 2024a); ChatTutor (Chen et al., 2024b); Park et al. (2024); Wang et al. (2025b); Liu et al. (2024c); Sonlu et al. (2024); Liu et al. (2024a) ,ss_work ] ] [Knowledge Tracing (KT), ss [Yang et al. (2024a); Xu et al. (2025); Scarlatos et al. (2025) ,ss_work ] ] [Error Correction and Detection (ECD), ss [Ye et al. (2022); Li et al. (2024c); Xu et al. (2024b); Repairagent (Bouzenia et al., 2024); Wang et al. (2025b); ErrorRadar (Yan et al., 2024c); CoT Rerailer (Wan et al., 2024) ,ss_work ] ] ] ] [Domain-Specific Educational Agents, ca [Science Learning §4.1, sl [Mathematics, sl [MathAgent (Yan et al., 2025a); TORA (Gou et al., 2023); MathChat (Wu et al., 2023); MACM (Lei et al., 2025); Xiong et al. (2024), sl_work ] ] [Physics, sl [NEWTON (Wang et al., 2023); Physics Reasoner (Pang et al., 2024); SGA (Ma et al., 2024c), sl_work ] ] [Chemistry, sl [ChemCrow (M. Bran et al., 2024); ChemAgent (Yu et al., 2024); ChemReasoner (Sprueill et al., 2024); Curie (Kon et al., 2025), sl_work ] ] [Biology, sl [ProtChat (Huang et al., 2024b); ProtAgents (Ghafarollahi and Buehler, 2024a); TourSynbio-Agent (Shen et al., 2024), sl_work ] ] [General Scientific Discovery, sl [PaperQA (Lála et al., 2023); PaperQA2 (Skarlinski et al., 2024); OpenScholar (Asai et al., 2024); SciAgents (Ghafarollahi and Buehler, 2024b); SciAgent (Ma et al., 2024a); SciToolAgent (Chen et al., 2025b); TheoremExplainAgent (Ku et al., 2025); LLM-SR (Shojaee et al., 2024); AI Scientist (Lu et al., 2024); ResearchAgent (Baek et al., 2024); Agent Laboratory (Schmidgall et al., 2025); DiscoveryWorld (Jansen et al., 2024), sl_work ] ] ] [Language Learning §4.2, ll [Reading, ll [ExpertEase (Mo and Hu, 2024); AgentSimp (Fang et al., 2025); Chen and Leitch (2024), ll_work ] ] [Writing, ll [Weaver (Wang et al., 2024b); CAELF (Hong et al., 2024); EvaAI (Lagakis and Demetriadis, 2024); Debate-to-Write (Hu et al., 2025a), ll_work ] ] [Translation, ll [Agent-SimT (Guo et al., 2024b); TransAgents (Wu et al., 2024), ll_work ] ], [Storytelling, ll [STARie (Li and Xu, 2023); StoryAgent (Sohn et al., 2024); LLaMS (Zhang et al., 2025c), ll_work ] ], [Speaking, ll [ELLMA-T (Pan et al., 2024); Spoken-WOZ (Si et al., 2023); FurChat (Cherakara et al., 2023), ll_work ] ] ] [Professional
Development §4.3, pd [Medical Education, pd [Talk2Care (Yang et al., 2024d); openCHA (Abbasian et al., 2023); MEDCO (Wei et al., 2024); Abd-Alrazaq et al. (2023); Steenstra et al. (2024); Agent Hospital (Li et al., 2024b), pd_work ] ] [Computer Science Education, pd [CodeAgent (Zhang et al., 2024b); ToolCoder (Ding et al., 2025); SWE-agent (Yang et al., 2025); AgentCoder (Huang et al., 2023c); MapCoder (Islam et al., 2024); Iris (Bassner et al., 2024); Chen et al. (2024d); HypoCompass (Ma et al., 2024b); AlgoBo (Jin et al., 2024a); Magentic-One (Fourney et al., 2024); AutoKaggle (Li et al., 2024e); AgentK v1.0 (Grosnit et al., 2024); SWE-Search (Antoniades et al., 2024); AgentLess (Xia et al., 2024); TRAVER (Wang et al., 2025a), pd_work ] ] [Law Education, pd [Lawbench (Fei et al., 2023); LegalBench (Guha et al., 2023); LegalAgentBench (Li et al., 2024a); DeliLaw (Xie et al., 2024a); Agentcourt (Chen et al., 2024c); Agentscourt (He et al., 2024a); AgentsBench (Jiang and Yang, 2024); LawLuo (Sun et al., 2024) , pd_work ] ] ] ] ]
Figure 2. Taxonomy of representative research on education agents.
3. Pedagogical Agents
Pedagogical agents are AI-driven systems designed to enhance both teaching and learning through automation and personalization. These agents can be broadly categorized into two types: Agents for Teaching Assistance, which automate key tasks for educators, and Agents for Student Support, which provide personalized learning guidance to students. By streamlining tasks for teachers and offering real-time, adaptive feedback to students, these agents improve instructional effectiveness, reduce workloads, and foster engaging, tailored learning experiences.
3.1. Agent for Teaching Assistance
@@
Agents for teaching assistance are designed to support educators in the learning environment. These agents leverage LLMs to provide personalized, scalable, and efficient support across various aspects of the educational process. Their primary objectives are to enhance teaching quality, enrich student learning experiences, and reduce educators’ workload. By incorporating advanced agent capabilities—such as memory management for retaining contextual information, tool integration for accessing external databases and APIs, and planning for precise student profile modeling—these agents can effectively assist in key areas, including classroom simulation3.1.1), feedback comment generation3.1.2), and learning resource recommendation3.1.3).
3.1.1. Classroom Simulation (CS)
@@
Effective classroom simulation depends on accurately modeling student behavior. Previous studies (Xu et al., 2024a; Jin et al., 2024b; Li et al., 2025a) demonstrate that LLM-based agents can predict fine-grained student behaviors across diverse personas and past learning patterns, aligning closely with real teachers’ expectations. To enhance simulation, the CGMI framework (Jinxin et al., 2023) uses a tree-based cognitive architecture with memory, reflection, and planning modules to simulate roles like teacher, student, and supervisor, improving realism. Similarly, Classroom Simulacra (Xu et al., 2025) incorporates a transferable iterative reflection module for more accurate behavior simulation. These systems enable automated interactions that reduce educators’ task loads while broadening the exploration of student profiles. Simulations can also test educational strategies tailored to different profiles, enhancing teaching quality, as shown in studies by Bhowmik et al. (2024) and Zheng et al. (2025). Additionally,  Hu et al. (2025b) demonstrate how LLMs can refine teaching plans through integrated simulations. To sum up, classroom simulations can be leveraged to test different educational strategies tailored to diverse student profiles, ultimately enhancing the quality of education.
3.1.2. Feedback Comment Generation (FCG)

### v2

@@
Abstract
Large Language Model (LLM) agents are transforming education by automating complex pedagogical tasks and enhancing both teaching and learning processes. In this survey, we present a systematic review of recent advances in applying LLM agents to address key challenges in educational settings, such as feedback comment generation, curriculum design, etc. We analyze the technologies enabling these agents, including representative datasets, benchmarks, and algorithmic frameworks. Additionally, we highlight key challenges in deploying LLM agents in educational settings, including ethical issues, hallucination and overreliance, and integration with existing educational ecosystems. Beyond the core technical focus, we include in Appendix A a comprehensive overview of domain-specific educational agents, covering areas such as science learning, language learning, and professional development.
LLM Agents for Education: Advances and Applications
Zhendong Chu1thanks: Equal contribution., Shen Wang [^1] footnotemark: 1, Jian Xie211footnotemark: 1, Tinghui Zhu211footnotemark: 1, Yibo Yan311footnotemark: 1, Jinheng Ye411footnotemark: 1, Aoxiao Zhong1, Xuming Hu3, Jing Liang1, Philip S. Yu5, Qingsong Wen1 1Squirrel Ai Learning, 2Fudan University 3Hong Kong University of Science and Technology (Guangzhou) 4Tsinghua University, 5University of Illinois Chicago {zc9uy@virginia.edu, qingsongedu@gmail.com}
1 Introduction
@@
Artificial intelligence techniques are increasingly used in education to enable personalized learning and intelligent tutoring Chen et al. (2020); Zhai et al. (2021); Pedro et al. (2019); Yan et al. (2025b). While traditional educational data mining approaches Shafiq et al. (2022); Khan and Ghosh (2021); Abdelrahman et al. (2023); Song et al. (2022); Wang et al. (2022); Gao et al. (2021), such as knowledge tracing and cognitive diagnosis, have made significant progress in reshaping the human-learning paradigm by analyzing student behaviors and assessing knowledge states, they still face major challenges in real-world applications. These challenges include shallow contextual understanding, limited interactive capabilities, and difficulties in generating adaptive, personalized learning materials, etc Zhu et al. (2024); Laak and Aru (2024); Tan et al. (2023).
@@
The strong natural language understanding of Large Language Models (LLMs) and the task automation capabilities of LLM agents make them valuable for addressing challenges in education Weng (2023); Wang et al. (2024a). First, memory enables LLM agents to retain both long-term knowledge about students’ study habits and short-term context from real-time interactions, enhancing contextual understanding and ensuring personalized learning experiences across various educational tasks Zhang et al. (2024d). Second, tool use allows LLM agents to access external resources, perform complex calculations, and retrieve real-time information, enabling them to automate intricate educational tasks such as grading, knowledge retrieval, and adaptive content generation, thereby overcoming limited interactivity and enhancing engagement Gao et al. (2023). Third, planning supports structured learning by decomposing complex topics, predicting optimal learning paths, and dynamically adjusting instructional strategies, allowing LLM agents to autonomously guide students through personalized learning experiences Huang et al. (2024b).
In addition to these core architectural capabilities, we identify personalization, explainability, and multi-agent communication as essential for effective educational LLM agents, as they enable nuanced instructional behavior and collaborative reasoning. Personalization allows agents to tailor instruction to individual learners’ needs and preferences Razafinirina et al. (2024). Explainability enables them to provide interpretable justifications for feedback and decisions Abu-Rasheed et al. (2024b). Multi-agent communication facilitates role-based collaboration, such as between a planner and a critic, to improve robustness and task coverage Wu et al. (2023a). By integrating these features, LLM agents enhance understanding and engagement while streamlining educational workflows for greater adaptability and efficiency.
In this survey, we provide a comprehensive review of LLM agents in educational settings, with a focus on their underlying technical foundations and general-purpose pedagogical capabilities. We begin by introducing the core abilities of educational LLM agents, including memory, tool use, planning, personalization, explainability, and multi-agent communication, and discuss their potential to automate and enhance diverse educational tasks. Given the highly applied nature of the education domain, we propose a task-centric taxonomy in Figure 1 that organizes recent advances around core educational tasks. We categorize educational LLM agents based on their roles in supporting teachers and students, capturing both instructional and learning-focused functions: (1) Teaching Assistance Agents, which support educators by automating tasks such as Classroom Simulation (CS), Feedback Comment Generation (FCG), and Curriculum Design (CD); and (2) Student Support Agents, which facilitate personalized learning through Adaptive Learning (AL), Knowledge Tracing (KT), and Error Correction and Detection (ECD). The overview of LLM agents for education and illustrative examples of each task are presented in Figure 2. We further highlight critical challenges in deploying these agents, including ethical issues, hallucination and overreliance, and integration into existing educational ecosystems. Finally, we compile essential datasets, benchmarks, and evaluation protocols to support future research on LLM agent-based educational systems. We summarize our contributions as follows:
  • Novel task-centric taxonomy. We propose a structured taxonomy that categorizes LLM agents based on core educational tasks, highlighting roles in teaching assistance and student support to unify analysis across applications.
@@
  • Current challenges and future directions. We analyze critical challenges that need to be addressed for the effective deployment of LLM agents for education, including issues related to ethical issues, hallucination, and integration into real-world educational ecosystems.
@@
  • Compilation of essential resources. We compile comprehensive datasets and benchmarks to support future research efforts and facilitate the development of more robust and effective LLM-driven educational solutions. [^1] Due to the page limit, we present more details in Appendix B
Beyond the main technical focus, Appendix A provides an overview of domain-specific educational agents, including applications in science learning, language learning, and professional development. We outline domain-specific challenges and review recent advances, benchmarks, and datasets. Readers are encouraged to refer to the appendix for further details.
@@
Figure 1: Taxonomy of representative research on LLM agents for education.
2 LLM Agents for Education
LLM agents have a set of connected abilities that help them handle complex tasks and provide meaningful support in education. These core features work together to let LLM agents do more than just find information—they can interact in ways that are flexible, responsive, and tailored to each learner. Their main strengths in education include strong memory, the ability to use tools, planning skills, personalization, clear explanations, and the ability to work with other agents. Specifically,
Memory. Memory in LLM agents includes long-term (foundational knowledge, e.g., commonsense) and short-term (current interaction data) components (Sumers et al., ). Active management via summarization and retrieval ensures relevant context is maintained (Chen et al., 2023; Liang et al., 2023; Zhong et al., 2023). This allows agents to track student progress and personalize responses, though sophisticated filtering is crucial to maintain the quality against noisy interaction data.
Tool Use. To overcome limitations like knowledge cutoffs or calculation difficulties, LLM agents use external tools such as search engines, databases, or APIs (Qin et al., 2023; Zhuang et al., 2023). This expands their functionality, providing access to current information and specialized capabilities. By integrating these tools, LLM agents expand their functional capabilities, ensuring that they can provide accurate and relevant information while supporting diverse educational tasks.
Planning. Planning allows agents to actively support learning by breaking down complex goals into smaller, manageable steps and adjusting based on how the student is doing (Yao et al., 2023; Valmeekam et al., 2023). This means understanding the learning goals, creating step-by-step plans, personalizing the learning path, and making changes based on feedback and progress. For longer-term tasks, LLMs can act as “meta-controllers,” using methods like “Pedagogical Steering” to stay aligned with teaching goals, which thus leads to dynamic, emergent curricula co-created with students (Zhang et al., 2025c).
Personalization. Personalization is a hallmark of effective education, and LLM agents excel in this area by adapting to individual learning styles and needs (Chen et al., ). They can serve as tutors, teaching assistants, or even simulate peer interactions to enhance the learning experience (Jin et al., 2024b; Guo et al., 2024c). Additionally, LLM agents can simulate user behaviors in recommendation systems, which can be applied to personalize educational content (Li et al., 2024c; Chu et al., 2025). Such a technology of personalization transforms traditional education, making it more accessible and tailored to each student’s unique needs.
Explainability. Explainability is crucial for building trust and facilitating learning in educational settings (Wu et al., 2024b). LLM agents must provide clear, step-by-step explanations that are understandable to students. Clear explanations are particularly important in subjects like STEM, where step-by-step reasoning is essential for student understanding (Nair et al., 2024; Zhang et al., 2025a). Consequently, the focused effort required to make educational LLM agents more transparent and their decision-making processes more scrutable.
Multi-Agent Communication. LLM agents can facilitate richer interactions in collaborative learning environments (Chen et al., ; Wu et al., 2023a). Multiple agents could coordinate on group projects or simulate peer learning, fostering critical thinking and diverse perspectives (Jinxin et al., 2023; Yue et al., 2024). This approach could support complex simulations of teamwork and real-world professional settings, especially when guided by new teaching frameworks for organization and evaluation (Xu et al., 2024a).
Table 1 summarizes the core capabilities required by each educational task supported by LLM agents. For each task, we highlight the primary capabilities identified from the reviewed literature. However, certain studies have also explored additional capabilities beyond these primary ones to enhance specific tasks—for example, Guo et al. (2024c) leverage multi-agent communication to improve feedback correctness. We omit these secondary capabilities from the table to more clearly illustrate the primary focus areas associated with each task.
Figure 2: The overview of LLM Agents for education. Teachers and students interact with LLM agents by submitting task-specific prompts. The agents respond using core capabilities such as memory, planning, tool use, and personalization to carry out tasks that support instruction and learning.
Task Memory Tool Use Planning Personalization Explainability Multi-Agent Comm.
CS
FCG
CD
AL
KT
ECD
Table 1: Mapping of educational tasks to essential LLM agent capabilities. $\checkmark$ indicates a primary capability; others may also be used as supporting components in some systems.
3 Agent for Teaching Assistance
@@
Agents for teaching assistance are designed to support educators in the learning environment. These agents leverage LLMs to provide personalized, scalable, and efficient support across various aspects of the educational process. Their primary objectives are to enhance teaching quality, enrich student learning experiences, and reduce educators’ workload. By incorporating advanced capabilities, LLM agents can effectively support key educational tasks, including classroom simulation (§3.1), feedback comment generation (§3.2), and curriculum design (§3.3).
3.1 Classroom Simulation
@@
As shown in Table 1, effective classroom simulation relies on memory, planning, and multi-agent communication to accurately model student behavior. Previous studies (Xu et al., 2024a; Jin et al., 2024b; Li et al., 2025a) demonstrate that LLM-based agents can predict fine-grained student behaviors across diverse personas and past learning patterns, aligning closely with real teachers’ expectations. To enhance simulation, the CGMI framework (Jinxin et al., 2023) uses a tree-based cognitive architecture with memory, reflection, and planning modules to simulate roles like teacher, student, and supervisor, improving realism. Similarly, Classroom Simulacra (Xu et al., 2025) incorporates a transferable iterative reflection module for more accurate behavior simulation. These systems enable automated interactions that reduce educators’ task loads while broadening the exploration of student profiles. Simulations can also test educational strategies tailored to different profiles, enhancing teaching quality, as shown in studies by Bhowmik et al. (2024) and Zheng et al. (2025a). Additionally,  Hu et al. (2025a) demonstrate how LLMs can refine teaching plans through integrated simulations. To sum up, classroom simulations can be leveraged to test different educational strategies tailored to diverse student profiles, ultimately enhancing the quality of education.
3.2 Feedback Comment Generation
Providing timely, relevant, and constructive feedback is a cornerstone of effective education. Teaching assistance agents can generate automated feedback comments on students’ assignments, quizzes, and projects.